OpenStudy (anonymous):

Solve 2|5x − 2| + 4 = 18 for x. Show your work to receive full credit.

OpenStudy (raden):

first, subtracted by 4 to both sides, giving us 2|5x − 2| + 4 - 4 = 18 - 4 2|5x − 2| = 14

OpenStudy (raden):

then divided by 2 to both sides : 2|5x − 2|/2 = 14/2 |5x − 2| = 7

OpenStudy (raden):

now, if given an equation |f(x)| = c, with c a constant then the solution satisfies when f(x) = c or -f(x) = c

OpenStudy (raden):

so, if given an equation |5x − 2| = 7, its solution is 5x − 2 = 7 or -(5x − 2) = 7 solve for x

OpenStudy (raden):

case I : 5x − 2 = 7 added by 2 to both sides : 5x -2 +2 = 7+2 5x = 9 divided by 5, 5x/5 = 9/5 x =9/5

OpenStudy (raden):

case II : -(5x − 2) = 7 multiply by -1 to both sides : -(5x − 2)(-1) = 7(-1) 5x − 2 = -7 added by 2 to both sides : 5x - 2 + 2 = -7 + 2 5x = -5 divided by 5 to both sides : 5x/5 = -5/5 x = -1

OpenStudy (raden):

so, the solution is {-1, 9/5}

OpenStudy (anonymous):

THANK YOU! THANK YOU! THANK YOU! I tried solving this problem & failed but thank you for helping me and using your time to help me with this!!! :)

OpenStudy (raden):

very welcome ::)